Determine value of (k) as an exact decimal so that the system has no solution.
5x-4y=8
6x+ky=11

Answer by stanbon(57320)

Determine value of (k) as an exact decimal so that the system has no solution.
5x-4y=8
6x+ky=11
--------------
The slope of the 1st equation is (5/4)
The slope of the 2nd equation is (-6/k)
----------------
To have no solution the slopes must te equal:
-6/k = 5/4
5k = -24
k = -24/5
---------------
If k=-24/5 the system will have no solution.
===============
cheers,
Stan H.